What is the magnitude of the acceleration of a spinster running at 10 m/s when Rounding A turn of radius 25 m?
1
Expert's answer
2019-12-16T11:10:50-0500

The magnitude of the acceleration of a spinster:


"a=\frac{v^2}{r}"

"a=\frac{10^2}{25}=4\frac{m}{s^2}"